OpenStudy (anonymous):

A pastry recipe calls for 3 cups of flour to make 8 servings. How many cups of flour are needed to make 6 servings? A. 2.25 cups B. 2.5 cups C. 4.5 cups D. 16 cups Twenty stamps cost $8.40. How much does it cost to buy 5 stamps? A. $42.00 B. $11.90 C. $4.20 D. $2.10

OpenStudy (anonymous):

3/8 = x/6 stamps are $8.40/20 , so 1 = ? so 5 = ?

OpenStudy (anonymous):

3/8 = x/6 x= 9/4

OpenStudy (anonymous):

Yes, 9/4 or 2 1/4 or 2.25 good.

OpenStudy (anonymous):

Then what do I do now?

OpenStudy (anonymous):

Chose answer A for your first problem. For your second problem $8.40/20 stamps is 0.42 per 1 stamp and 5 x 0.42 = $2.10 for 5 and you are done.

OpenStudy (anonymous):

Okay thanks so much!!!